Can You plesae help me.
3/4=1-3x-2/x+1
This is what I got
3/4=3-2/1
3/4=1/1
3/4=1
Thank you

Answer by BurningFlame_marc(77)

Hello!
There's something wrong in your solution. First, you shouldn't eliminate the variable because your objective is to find the value of the variable.

Now here's my solution.
3/4=1-3x-2/x+1
3/4(x+1)=1-3x-2
3/4x+3/4=-3x-1
3/4x=-3x-7/4
15/4x=-7/4
15x=-7
x=-7/15
The answer is -7/15. If you have any questions, it will be my honor to teach you.